OpenStudy (anonymous):

I need to identify the conic section for these 10 questions. I need help to find out whether they are Elipse, Hyperbole, or Circle. Hold on I shall post the questions.

OpenStudy (anonymous):

1. x^2/9 + y^2/100 =1

OpenStudy (anonymous):

2. y^2/25 - x^2/1 = 1

OpenStudy (anonymous):

3.( x^2)/(28) - y = 0 4. (y^2)/(16) + (x^2) / (16) = 1

OpenStudy (anonymous):

5. (x^2) -( (y-8)^2 / (36) ) =1

OpenStudy (anonymous):

6. (1)/(80) y^2 =x 7. (y+2)^2 / (49) + (x^2)/ (4) = 1 8. (x-2)^2 + (y+3)^2 = 16

OpenStudy (anonymous):

9. x^2 -10x+25 = y 10. ((x-11)^2 / (144)) - ((y-1)^2 /(81)) =1

OpenStudy (anonymous):

@campbell_st @rebeccaskell94

OpenStudy (anonymous):

I think #1 is elipse and #2 is hyperbole

OpenStudy (anonymous):

and I think #10 is Elipse

OpenStudy (campbell_st):

here is a website I just googles that has a nice summary of conic section equations http://math2.org/math/algebra/conics.htm and 1 is an ellipse 2 is a hyperbola good luck

OpenStudy (anonymous):

wait am I right for #10?

OpenStudy (anonymous):

can someone please tell me 3-9? :(

OpenStudy (anonymous):

@mustry ?

OpenStudy (anonymous):

@hba ?

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

I need 3-10

OpenStudy (radar):

Wouldn't 4. be a circle?

OpenStudy (anonymous):

that's what I was thinking. what about 3?

OpenStudy (radar):

Ima lookin at 3.

OpenStudy (anonymous):

OH WAIT! it says Parabola is an option too

OpenStudy (anonymous):

Is 6 a parabola?

OpenStudy (radar):

Wouldn't 3 be a parabola??

OpenStudy (anonymous):

I believe so, what about 6?

OpenStudy (anonymous):

And isn't 10 ellipse?

OpenStudy (radar):

I think 6 also a parabola.

OpenStudy (anonymous):

I think so too. Is 10 ellipse?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@mustry

OpenStudy (radar):

I am not sure of 10. The fact that it is a difference confuses me otherwise it would be a circle (I think), but with the minus, I am confused.

OpenStudy (anonymous):

Hm, well what about 5, 7, 8, and 9?

OpenStudy (radar):

I think 9 is a parabola.

OpenStudy (radar):

8. would be a circle whose origin is not at (0,0)

OpenStudy (anonymous):

alright, I just need 5, 7, and 10

OpenStudy (anonymous):

isn't 7 a hyperbola?

OpenStudy (radar):

5 is similar to 10 (whatever that is !!)

OpenStudy (anonymous):

isn't 7 a hyperbola?

OpenStudy (radar):

7 circle and 5 and 10 hyperbola. These are guesses.

OpenStudy (radar):

I am sorry it is obvious I need to review the conics. lol

OpenStudy (anonymous):

thank you so much lol